OpenStudy (anonymous):

If possible, find the exact value of sec[arctan(-5)]. Pls explain steps to solve such questions....

OpenStudy (calculusfunctions):

arctan means inverse of the tangent function.\[\tan ^{-1}\]is another way of writing arctan (inverse of tan). If\[arc \tan y =\theta \]for example, then\[y =\tan \theta \]Do you understand? If this clue helps, then what do you think is the first step towards your problem?

OpenStudy (anonymous):

that way I will get -5 = tan A (i am using A for ease of typing here)...then what???? Do I draw a right triangle??? In which quadrant???

OpenStudy (anonymous):

If possible pls explain this question with full step by step solution so that I can go through it and get to know how to solve such sums...This is one of the few I could not do....

OpenStudy (calculusfunctions):

Example: Find the exact value of\[\sin (arc \cos 0.8)\] Solution: First let arc cos 0.8 = θ, this implies that cos θ = 0.8 = 4/5. Now sin²θ + cos²θ = 1 sin²θ + (4/5)² = 1 [∵ cos θ = 4/5] sin²θ + (16/25) = 1 sin²θ = 1 - (16/25) sin²θ = 9/25 sin θ = ± 3/5 We reject -3/5 because cos θ was positive. sin θ = 3/5 ∴ sin (arc cos 0.8) = 3/5

OpenStudy (calculusfunctions):

You could have also drawn a triangle.  From which you see that sin θ = 3/5. Thus you may use either the identities or the triangle.